Monthly Archives: October 2011

I500 I400 VXR400 Normal Service Restored

ITG has disabled their misconfigured backup process until it can be fixed. So long term runs should be reliable again.

Posted in Announcements | Comments Off

I500 I400 and VXR400 Operational Again

These instruments have been restarted again with PC reboots.

Posted in Announcements | Comments Off

500 and 400s working

The I500, I400 and VXR400 stopped working since last night due to a communication issue between the spectrometers and their host computers. They have been reset this morning and are working now.

Posted in Announcements | Comments Off

VXR400 sample ejection problem fixed

Last night, an user put a bare spinner (without a sample tube on it) into the magnet after his/her use of the spectrometer, which apparently causes problems as the spinner cannot be ejected out afterward. PLEASE DO NOT LET THIS HAPPEN AGAIN.

Posted in Announcements | Comments Off